5.3307 g of Na2CrO4 (MW 161.97 g/mol) is dissolved in 1000.0 mL of water. Assuming the solution has a density of 1.00 g/mL, what is the concentration of Na (MW 22.9898 g/mol) in the solution in units of a) molarity (M)? Number b) parts per thousand (ppt)? Number ppt c) 25.0 mL of the solution is then diluted to a final volume of 500.0 mL. What is the concentration of Na* in the diluted solution in units of parts per mllion (ppm)? Number Ppm
Answer #1

ans)

a)

Molarity

Molarity = mass / molecular wt X volume ( L)

Molarity Na2CrO4 = 5.3307 / 161.97 X 0.1000 = 0.329 M

So molarity of sodium = 2 X molarity of salt = 0.658 M

b) Parts per thousand (ppt)

parts / thousand = Molarity X mol wt = 0.658 X 23 = 15.134 g / L

c)

   25.0 mL of the solution is then diluted to a final volume of 500.0 mL. What is the concentration of Na in the diluted solution in units of parts per million (ppm)?

M1V1 = M2V2

25 X 0.658 = M2 X 500

M2 = 0.0329

Strength = Molarity X mol wt X 10^3 ppm = 0.7567 X 10^3 ppm